OpenStudy (anonymous):

Find the distance between the points (-2, 4) and (5, 4).

OpenStudy (anonymous):

6 , -1

OpenStudy (anonymous):

its not a distance

OpenStudy (michele_laino):

hint: we have to apply this formula: \[d = \sqrt {{{\left( {{x_2} - {x_1}} \right)}^2} + {{\left( {{y_2} - {y_1}} \right)}^2}} \] where (x_1,y_1) and (x_2,y_2) are your points

OpenStudy (anonymous):

\[\sqrt{(5-(-2))^2 +(4-4)^2}\]

OpenStudy (anonymous):

the answer choices are 3, 7, or 8.